OpenStudy (goformit100):

(x^4 + 2 x i) -(3x^2+ iy) = (3-5i)+(1+2iy) find x and y. a) x = 2, y = 3 b) x = 3, y = 1/3 c) x = -2, y = -3 d) none of these ((Show full working))

OpenStudy (gorv):

add similar term

OpenStudy (gorv):

x^4+2xi-3x^2-iy

OpenStudy (gorv):

(x^4-3x^2)+i(2x-y)

OpenStudy (gorv):

in complex when number are compared than real part is compared with real and i with i

OpenStudy (gorv):

(x^4-3x^2)+i(2x-y)= (3-5i)+(1+2iy) (x^4-3x^2)+i(2x-y)=3+1-5i+2iy (x^4-3x^2)+i(2x-y)=4+i(-5+2y)

OpenStudy (gorv):

compare real with real and i with i now

OpenStudy (gorv):

x^4-3x^2=4

OpenStudy (gorv):

and 2x-y=-5+2y

OpenStudy (gorv):

solve these two to get your ans

OpenStudy (gorv):

its simple to solve now u can do it by yourself

OpenStudy (goformit100):

ok how to solve further ?

OpenStudy (gorv):

2x-y=-5+2y 2x-3y=-5

OpenStudy (gorv):

x^4-3x^2=4 factorise this one

OpenStudy (gorv):

and put value of x from here in 2x-3y=-5

OpenStudy (gorv):

to get y

OpenStudy (goformit100):

ok

OpenStudy (gorv):

x^4-3x^2=4 hint: let x^2=t

OpenStudy (gorv):

t^2-3t=4

OpenStudy (gorv):

t^2-3t-4=0

OpenStudy (gorv):

t^2-4t+t-4=0 t(t-4)+(t-4)=0 (t+1)(t-4)=0

OpenStudy (gorv):

and remember t=x^2 so whatever the value of t put it =x^2

OpenStudy (gorv):

t=-1 and t=4

OpenStudy (gorv):

x^2=-1 not possible it wont give you the negative value ...so cancel this solution

OpenStudy (gorv):

x^2=4

OpenStudy (gorv):

x=-2 or 2

OpenStudy (gorv):

now try further plzzz

OpenStudy (goformit100):

thak you

OpenStudy (gorv):

welcm buddy :)